Find the integral of (y² -1)dx -2dy =0
1
Expert's answer
2020-12-01T01:38:22-0500

"\displaystyle


 (y² -1)\mathrm{d}x -2\mathrm{d}y =0\\

y^2 - 1 - 2y' = 0\\

y^2 - 1 = 2y' \\



\frac{2}{y^2 - 1} \,\mathrm{d}y = \mathrm{d}x\\

\int\frac{2}{y^2 - 1} \,\mathrm{d}y = \int\mathrm{d}x\\


\frac{1}{2}\cdot 2 \ln\left(\frac{y - 1}{y + 1}\right) = x + C\\

\frac{y - 1}{y + 1} = e^{x + C}\\


\frac{y - 1}{y + 1} = Ae^{x}\\


y - 1 = A(y + 1)e^{x}\\


y(1 - Ae^{x}) = Ae^{x} + 1 \\


y = \frac{1 + Ae^{x}}{1 - Ae^{x}}"